### v1

# Pick up the Pace: A Parameter-Free Optimizer for Lifelong Reinforcement Learning

Abstract
A key challenge in lifelong reinforcement learning (RL) is the loss of plasticity, where previous learning progress hinders an agent’s adaptation to new tasks. While regularization and resetting can help, they require precise hyperparameter selection at the outset and environment-dependent adjustments. Building on the principled theory of online convex optimization, we present a parameter-free optimizer for lifelong RL, called Pace, which requires no tuning or prior knowledge about the distribution shifts. Extensive experiments on Procgen, Atari, and Gym Control environments show that Pace works surprisingly well—mitigating loss of plasticity and rapidly adapting to challenging distribution shifts—despite the underlying optimization problem being nonconvex and nonstationary.
1 Introduction
Figure 1: Severe loss of plasticity in Procgen (Starpilot). There is a steady decline in reward with each distribution shift.
Spot, the agile robot dog, has been learning to walk confidently across soft, lush grass. But when Spot moves from the grassy field to a gravel surface, the small stones shift beneath her feet, causing her to stumble. When Spot tries to walk across a sandy beach or on ice, the challenges multiply, and her once-steady walk becomes erratic. Spot wants to adjust quickly to these new terrains, but the patterns she learned on grass are not suited to gravel, sand, or ice. Furthermore, she never knows when the terrain will change again and how different it will be, therefore must continually plan for the unknown while avoiding reliance on outdated experiences.
Spot’s struggle exemplifies a critical challenge in real-world decision making, known as lifelong reinforcement learning (lifelong RL). The idea is that the learning agent must continually acquire new knowledge to handle the nonstationarity of the environment. At first glance, there appears to be an obvious solution: given a policy gradient oracle, the agent could just keep running gradient descent nonstop. However, recent experiments have demonstrated an intriguing behavior called loss of plasticity (Dohare et al., 2021; Lyle et al., 2022; Abbas et al., 2023; Sokar et al., 2023): despite persistent gradient steps, such an agent can gradually lose its responsiveness to incoming observations. There are even extreme cases of loss of plasticity (known as negative transfer or primacy bias), where prior learning can significantly hamper the performance in new tasks (Nikishin et al., 2022; Ahn et al., 2024); see Figure 1 for an example. All these suggest that the problem is more involved than one might think.
From the optimization perspective, the above issues might be attributed to the lack of stability under gradient descent. That is, the weights of the agent’s parameterized policy can drift far away from the origin (or a good initialization), leading to a variety of undesirable behaviors. [^1] Such as the inactivation of many neurons, due to the ReLU activation function (Abbas et al., 2023; Sokar et al., 2023). Fitting this narrative, it has been shown that simply adding a $L_{2}$ regularizer to the optimization objective (Kumar et al., 2023) or periodically resetting the weights (Dohare et al., 2021; Asadi et al., 2023; Sokar et al., 2023; Ahn et al., 2024) can help mitigate the problem. However, a particularly important limitation is their use of hyperparameters, such as the magnitude of the regularizer and the resetting frequency. Good performance hinges on the suitable environment-dependent hyperparameter, but how can one confidently choose that before interacting with the environment? The classical cross-validation approach would violate the one-shot nature of lifelong RL (and online learning in general; see Chapter 1 of Orabona, 2023), since it is impossible to experience the same environment multiple times. This leads to the contributions of the present work.
Contribution
The present work addresses the key challenges in lifelong RL using the principled theory of Online Convex Optimization (OCO). Specifically, our contributions are two fold.
  • Algorithm: Pace Building on a series of results in OCO (Cutkosky & Orabona, 2018; Cutkosky, 2019; Cutkosky et al., 2023; Zhang et al., 2024b), we propose a (hyper)-parameter-free optimizer for lifelong RL, called Pace (Parameter-free Adaption for Continual Environments). Intuitively, the idea is a refinement of regularization: instead of manually selecting the magnitude of regularization beforehand, Pace chooses that in an online, data-dependent manner. From the perspective of OCO theory, Pace is insensitive to its own hyperparameter, which means that no hyperparameter tuning is necessary in practice. Furthermore, as an optimization approach to lifelong RL, Pace is compatible with any policy parameterization.
  • Experiment Using Proximal Policy Optimization (PPO), we conduct comprehensive experiments on the instantiation of Pace called Pace PPO. A diverse range of lifelong RL environments are tested (based on Procgen, Atari, and Gym Control), with considerably larger scale than prior works. In settings where existing approaches (Abbas et al., 2023; Kumar et al., 2023) struggle, we find that Pace PPO
    • mitigates mild and extreme loss of plasticity;
    • and rapidly adapts to new tasks when distribution shifts are introduced.
    Such findings might be surprising: the theoretical advantage of Pace is motivated by the convexity in OCO, but lifelong RL is both nonconvex and nonstationary in terms of optimization.
Organization
Section 2 surveys the basics of lifelong RL. Section 3 introduces our parameter-free algorithm Pace, and experiments are presented in Section 4. We defer the discussion of related works and results to Section 5. Finally, Section 6 concludes the paper.
2 Lifelong RL
As a sequential decision making framework, reinforcement learning (RL) is commonly framed as a Markov Decision Process (MDP) defined by the state space ${\cal S}$, the action space ${\cal A}$, the transition dynamics $P(s_{t+1}|s_{t},a_{t})$, and the reward function $R(s_{t},a_{t},s_{t+1})$. In the $t$-th round, starting from a state $s_{t}\in{\cal S}$, the learning agent needs to choose an action $a_{t}\in{\cal A}$ without knowing $P$ and $R$. Then, the environment samples a new state $s_{t+1}\sim P(\cdot|s_{t},a_{t})$, and the agent receives a reward $r_{t}=R(s_{t},a_{t},s_{t+1})$. There are standard MDP objectives driven by theoretical tractability, but from a practical perspective, we measure the agent’s performance by its cumulative reward $\sum_{t=1}^{T}r_{t}$.
The standard setting above concerns a stationary MDP. Motivated by the prevalence of distribution shifts in practice, the present work studies a nonstationary variant called lifelong RL, where the transition dynamics $P_{t}$ and the reward function $R_{t}$ can vary over time. Certainly, one should not expect any meaningful “learning” against arbitrary unstructured nonstationarity. Therefore, we implicitly assume $P_{t}$ and $R_{t}$ to be piecewise constant over time, and each piece is called a task – just like our example of Spot in the introduction. The main challenge here is to transfer previous learning progress to new tasks. This is reasonable when tasks are similar, but we also want to reduce the degradation when tasks turn out to be very different.
Lifelong RL as online optimization
Modern RL approaches, including PPO (Schulman et al., 2017) and others, crucially utilize the idea of policy parameterization. Specifically, a policy refers to the distribution of the agent’s action $a_{t}$ (conditioned on the historical observations), and we use $\theta_{t}\in\mathbb{R}^{d}$ to denote the parameterizing weight vector. After sampling $a_{t}$ and receiving new observations, the agent could define a loss function $J_{t}(\theta)$ that characterizes the “hypothetical performance” of each weight $\theta\in\mathbb{R}^{d}$. Then, by computing the policy gradient $g_{t}=\nabla J_{t}(\theta_{t})$, one could apply a first order optimization algorithm [^2] Formally, a dynamical system that given its state $\theta_{t}$ and input $g_{t}$ outputs the new state $\textsc{OPT}(\theta_{t},g_{t})$. OPT to obtain the updated weight, $\theta_{t+1}=\textsc{OPT}(\theta_{t},g_{t})$.
For the rest of this paper, we will work with such an abstraction. The feedback of the environment is treated as a policy gradient oracle ${\cal G}$, which maps the time $t$ and the current weight $\theta_{t}$ into a policy gradient $g_{t}={\cal G}(t,\theta_{t})$. Our goal is to design an optimizer OPT well suited for lifelong RL.
Lifelong vs. Continual
In the RL literature, the use of “lifelong” and “continual” varies significantly across studies, which may lead to confusion. Abel et al. (2023) characterized continual reinforcement learning (CRL) as a never-ending learning process. However, much of the literature cited under CRL, such as (Abbas et al., 2023; Ahn et al., 2024), primarily focuses on the problem of backward transfer (avoiding catastrophic forgetting). Conversely, the present work addresses the problem of forward transfer, which refers to the rapid adaptation to new tasks. Because of this we use “lifelong” rather than “continual” in our exposition, similar to (Thrun, 1996; Abel et al., 2018b; Julian et al., 2020).
3 Method
Figure 2: Visualization of Pace’s key idea.
Inspired by (Cutkosky et al., 2023), we study lifelong RL by exploiting its connection to Online Convex Optimization (OCO; Zinkevich, 2003). The latter is a classical theoretical problem in online learning, and much effort has been devoted to designing parameter-free algorithms that require minimum tuning or prior knowledge (Streeter & Mcmahan, 2012; McMahan & Orabona, 2014; Orabona & Pál, 2016; Foster et al., 2017; Cutkosky & Orabona, 2018; Mhammedi & Koolen, 2020; Chen et al., 2021; Jacobsen & Cutkosky, 2022). The surprising observation of Cutkosky et al. (2023) is that several algorithmic ideas closely tied to the convexity of OCO can actually improve the nonconvex deep learning training, suggesting certain notions of “near convexity” on its loss landscape. We find that lifelong RL (which is both nonconvex and nonstationary in terms of optimization) exhibits a similar behavior, therefore a particularly strong algorithm (named Pace) can be obtained from principled results in parameter-free OCO. Let us start from the background.
Basics of (parameter-free) OCO
As a standalone theoretical topic, OCO concerns a sequential optimization problem where the convex loss function $l_{t}$ can vary arbitrarily over time. In the $t$-th iteration, the optimization algorithm picks an iterate $x_{t}$ and then observes a gradient $g_{t}=\nabla l_{t}(x_{t})$. Motivated by the pursuit of “convergence” in optimization, the standard objective is to guarantee low (i.e., sublinear in $T$) static regret, defined as
$\mathrm{Regret}_{T}(l_{1:T},u)\mathrel{\mathop{:}}=\sum_{t=1}^{T}l_{t}(x_{t})-\sum_{t=1}^{T}l_{t}(u),$
where $T$ is the total number of rounds, and $u$ is a comparator that the algorithm does not know beforehand. In other words, the goal is to make $\mathrm{Regret}_{T}(l_{1:T},u)$ small for all possible loss sequence $l_{1:T}$ and comparator $u$. Note that for nonstationary OCO problems analogous to lifelong RL, it is better to consider a different objective called the discounted regret. Algorithms there mostly follow the same principle as in the stationary setting, just wrapped by loss rescaling (Zhang et al., 2024a).
For minimizing static regret, classical minimax algorithms like gradient descent (Zinkevich, 2003) would assume a small uncertainty set ${\cal U}$ at the beginning. Then, by setting the hyperparameter (such as the learning rate) according to ${\cal U}$, it is possible to guarantee sublinear worst case regret,
$\max_{(l_{1:T},u)\in{\cal U}}\mathrm{Regret}_{T}(l_{1:T},u)=o(T).$ (1)
In contrast, parameter-free algorithms use very different strategies [^3] The key difference with gradient descent is the use of intricate (non-$L_{2}$) regularizers. See (Fang et al., 2022; Jacobsen & Cutkosky, 2022) for a theoretical justification of their importance. to bound $\mathrm{Regret}_{T}(l_{1:T},u)$ directly (without taking the maximum) by a function of both $l_{1:T}$ and $u$. The resulting bound is more refined than Eq.(1) (Orabona, 2023, Chapter 9), and crucially, since there is no need to pick an uncertainty set ${\cal U}$, much less hyperparameter tuning is needed. This is where its name comes from.
Pace for lifelong RL
Now back to lifelong RL. As we discussed, a fundamental challenge here is the excessive drifting of the weights $\theta_{t}$, and this can be fixed by enforcing the proximity to a good reference point $\theta_{\mathrm{ref}}$. Different from existing approaches like $L_{2}$ regularization and resetting, parameter-free OCO provides a principled solution to this problem without hyperparameter-tuning. Naming this algorithm as Pace, we present its generic template as Algorithm 1, which calls Algorithm 2 (a one-dimensional scale tuner) as the key subroutine.
From the technical perspective, Pace assembles three techniques in parameter-free OCO: the direction-magnitude decomposition from (Cutkosky & Orabona, 2018), the additive aggregation from (Cutkosky, 2019), and the $\mathrm{erfi}$ potential function from (Zhang et al., 2024b). In addition, we follow (Cutkosky et al., 2023) for the recommended default parameter setup.
Algorithm 1 Pace: Parameter-free Adaption for Continual Environments.
1:  Input: A policy gradient oracle ${\cal G}$; a first order optimization algorithm Base; a reference point $\theta_{\mathrm{ref}}\in\mathbb{R}^{d}$; $n$ discount factors $\beta_{1},\ldots,\beta_{n}\in(0,1]$ (default: $0.9,0.99,\ldots,0.999999$).
2:  Initialize: Create $n$ copies of Algorithm 2, denoted as ${\cal A}_{1},\ldots,{\cal A}_{n}$. For each $j\in[1:n]$, ${\cal A}_{j}$ uses the discount factor $\beta_{j}$. Initialize the algorithm Base at $\theta_{\mathrm{ref}}$. Let $\theta_{1}=\theta_{\mathrm{ref}}$.
3:  for $t=1,2,\ldots$ do
4:     Obtain the $t$-th policy gradient $g_{t}={\cal G}(t,\theta_{t})\in\mathbb{R}^{d}$.
5:     Send $g_{t}$ to Base as its $t$-th input, and get its output $\theta_{t+1}^{\mathrm{Base}}\in\mathbb{R}^{d}$.
6:     For all $j\in[1:n]$, send $\left\langle g_{t},\theta_{t}-\theta_{\mathrm{ref}}\right\rangle$ to ${\cal A}_{j}$ as its $t$-th input, and get its output $s_{t+1,j}\in\mathbb{R}$.
7:     Define the scaling parameter $S_{t+1}=\sum_{j=1}^{n}s_{t+1,j}$.
8:     Update the weight of the policy,
$\theta_{t+1}=\theta_{\mathrm{ref}}+\left({\theta^{\mathrm{Base}}_{t+1}-\theta_{\mathrm{ref}}}\right)S_{t+1}.$
9:  end for
Algorithm 2 1D Discounted Tuner of Pace.
1:  Input: Discount factor $\beta\in(0,1]$; small value $\varepsilon>0$ (default: $10^{-8}$).
2:  Initialize: The running variance $v_{0}=0$; the running (negative) sum $\sigma_{0}=0$.
3:  for $t=1,2,\ldots$ do
4:     Obtain the $t$-th input $h_{t}$.
5:     Let $v_{t}=\beta^{2}v_{t-1}+h^{2}_{t}$, and $\sigma_{t}=\beta\sigma_{t-1}-h_{t}$.
6:     Select the $t$-th output
$s_{t+1}=\frac{\varepsilon}{\mathrm{erfi}(1/\sqrt{2})}\mathrm{erfi}\left({\frac{\sigma_{t}}{\sqrt{2v_{t}}+\varepsilon}}\right),$
where $\mathrm{erfi}$ is the imaginary error function queried from standard software packages.
7:  end for
Without going deep into the theory, here is an overview of the important ideas (also see Figure 2 for a visualization).
  • First, Pace is a meta-algorithm that operates on top of a “default” optimizer Base. It can simply be gradient descent with a constant learning rate, or Adam (Kingma & Ba, 2014) as in our experiments. Applying Base alone would be equivalent to enforcing the scaling parameter $S_{t+1}\equiv 1$ in Pace, but this would suffer from the drifting of $\theta_{t+1}^{\mathrm{Base}}$ (and thus, the weight $\theta_{t+1}$).
  • To fix this issue, Pace uses the tuner (Algorithm 2) to select the scaling parameter $S_{t+1}$, making it data-dependent. Typically $S_{t+1}$ is within $[0,1]$ (see Figure 12 to 14), therefore essentially, we define the updated weight $\theta_{t+1}$ as a convex combination of the Base’s weight $\theta_{t}^{\mathrm{Base}}$ and the reference point $\theta_{\mathrm{ref}}$,
    $\theta_{t+1}=S_{t+1}\cdot\theta_{t+1}^{\mathrm{Base}}+(1-S_{t+1})\theta_{\mathrm{ref}}.$
    This brings the weight closer to $\theta_{\mathrm{ref}}$, which is known to be “safe” (i.e., not overfitting any particular lifelong RL task), although possibly conservative.
  • To inject the right amount of conservatism without hyperparameter tuning, the tuner (Algorithm 2) applies an unusual decision rule based on the $\mathrm{erfi}$ function. Theoretically, this is known to be optimal in an idealized variant of OCO (Zhang et al., 2022, 2024b), but removing the idealized assumptions requires a tiny bit of extra conservatism, which is challenging (and not necessarily practical). Focusing on the lifelong RL problem that considerably deviates from OCO, we simply apply the $\mathrm{erfi}$ decision rule as is. This is loosely motivated by deep learning training dynamics, e.g., (Cohen et al., 2020; Ahn et al., 2023; Andriushchenko et al., 2023), where an aggressive optimizer is often observed to be better.
  • Finally, the tuner requires a discount factor $\beta$. This crucially controls the strength of regularization (elaborated next), but also introduces a hyperparameter tuning problem. Following (Cutkosky, 2019), we aggregate tuners with different $\beta$ (on a log-scaled grid) by simply summing up their outputs. This is justified by the adaptivity of the tuner itself: in OCO, if we add a parameter-free algorithm ${\cal A}_{1}$ to any other algorithm ${\cal A}_{2}$ that already works well, then ${\cal A}_{1}$ can automatically identify this and “tune down” its aggressiveness, such that ${\cal A}_{1}+{\cal A}_{2}$ still performs as well as ${\cal A}_{2}$.
Connection to regularization
Despite its nested structure, Pace can actually be seen as a parameter-free refinement of $L_{2}$ regularization (Kumar et al., 2023). To concretely explain this intuition, let us consider the following two optimization dynamics.
  • First, suppose we run gradient descent with learning rate $\eta$, on the policy gradient sequence $\{g_{t}\}$ with the $L_{2}$ regularizer $\frac{\lambda}{2}\left\|\theta-\theta_{\mathrm{ref}}\right\|^{2}$. Quantitatively, it means that starting from the $t$-th weight $\theta_{t}$,
    $\theta_{t+1}=\theta_{t}-\eta\left[{g_{t}+\lambda\left({\theta_{t}-\theta_{\mathrm{ref}}}\right)}\right],\quad\Longrightarrow\quad\theta_{t+1}-\theta_{\mathrm{ref}}=\left({1-\lambda\eta}\right)\left({\theta_{t}-\theta_{\mathrm{ref}}}\right)-\eta g_{t}.$ (2)
    That is, the updated weight $\theta_{t+1}$ is determined by a ($1-\lambda\eta$)-discounting with respect to the reference point $\theta_{\mathrm{ref}}$, followed by a gradient step $-\eta g_{t}$.
  • Alternatively, consider applying the following simplification of Pace on the same policy gradient sequence $\{g_{t}\}$: ($i$) Base is still gradient descent with learning rate $\eta$; ($ii$) there is just one discount factor $\beta$; and ($iii$) the one-dimensional tuner (Algorithm 2) is replaced by the $\beta$-discounted gradient descent with learning rate $\alpha$, i.e., $S_{t+1}=\beta S_{t}-\alpha h_{t}$. In this case, we have
    $\theta_{t+1}-\theta_{\mathrm{ref}}$ $=S_{t+1}\left({\theta^{\mathrm{Base}}_{t+1}-\theta_{\mathrm{ref}}}\right)$
    $=\left({\beta S_{t}-\alpha h_{t}}\right)\left({\theta^{\mathrm{Base}}_{t}-\theta_{\mathrm{ref}}-\eta g_{t}}\right)$
    $=\left({\beta-\alpha S^{-1}_{t}h_{t}}\right)\left({\theta_{t}-\theta_{\mathrm{ref}}}\right)-\eta S_{t+1}g_{t}.$ (mildly assuming $S_{t}\neq 0$)
    Notice that $S_{t}$ is a $\beta$-discounted sum of $\alpha h_{1},\ldots,\alpha h_{t-1}$, thus in the typical situation of $\beta\approx 1$ one might expect $\alpha h_{t}\ll|S_{t}|$. Then, the resulting update of $\theta_{t+1}$ is similar to Eq.(2), with quantitative changes on the “effective discounting” $1-\lambda\eta\rightarrow\beta$, and the “effective learning rate” $\eta\rightarrow\eta S_{t+1}$.
The main message here is that under a simplified setting, Pace is almost equivalent to $L_{2}$ regularization. The latter requires choosing the hyperparameters $\lambda$ and $\eta$, and similarly, the above simplified Pace requires choosing $\beta$ and $\eta$. Going beyond this simplification, the actual Pace removes the tuning of $\beta$ using aggregation, and the tuning of $\eta$ using the $\mathrm{erfi}$ decision rule.
On the hyperparameters
Although Pace is called “parameter-free”, it still needs the $\beta$-grid, the constant $\varepsilon$ and the algorithm Base as inputs. The idea is that Pace is particularly insensitive to such choices, as supported by the OCO theory. As the result, the generic default values recommended by Cutkosky et al. (2023) are sufficient in practice. We note that those are proposed for training supervised deep learning models, thus should be agnostic to the lifelong RL applications we consider.
4 Experiment
Figure 3: Experimental setup for lifelong RL.
Table 1: Cumulative sum of mean episode reward for Pace PPO, Adam PPO, and CReLU on Procgen, Atari, and Gym Control environments. Rewards are scaled by $10^{5}$; higher is better.
Environment Adam PPO CReLU Pace PPO (Ours)
Starpilot $3.4$ $3.6$ $\mathbf{12.5}$
Dodgeball $1.9$ $2.3$ $\mathbf{5.2}$
Chaser $1.4$ $1.7$ $\mathbf{2.2}$
Fruitbot $0.1$ $1.0$ $\mathbf{1.8}$
CartPole $5.1$ $1.2$ $\mathbf{39.6}$
Acrobot $-14.3$ $-13.9$ $\mathbf{-12.9}$
LunarLander $-21.7$ $-19.4$ $\mathbf{-8.6}$
Atari 6 $3.1$ $4.8$ $\mathbf{10.5}$
Atari 9 $3.9$ $17.0$ $\mathbf{20.2}$
Does Pace experience the common pitfalls of loss of plasticity? Does it rapidly adapt to distribution shifts? To answer these questions, we test Pace in empirical RL benchmarks such as vision-based games and physics-based control environments in lifelong settings (Figure 3). Specifically, we instantiate PPO with two different optimizers: Adam with constant learning rate for baseline comparison, and Pace for our proposed method (with exactly the same Adam as the input Base). We also test Adam PPO with concatenated ReLU activations (CReLU; Shang et al., 2016), previously shown to mitigate loss of plasticity in certain deep RL settings (Abbas et al., 2023). Our numerical results are summarized in Table 1. Across every lifelong RL setting, we observe substantial improvements in the cumulative episode reward by using Pace PPO compared to Adam PPO or CReLU. Below are the details, with more in Appendix E.
OpenAI Procgen
We first evaluate on OpenAI Procgen, a suite of 16 procedurally generated game environments (Cobbe et al., 2019). We introduce distribution shifts by sampling a new procedurally generated level of the current game every 2 million time steps, treating each level as a distinct task.
Figure 4: Reward in the lifelong Procgen environments for StarPilot, Dodgeball, Fruitbot, and Chaser. The plots demonstrate a steady loss of plasticity in agents using Adam PPO and CReLU, characterized by their inability to maintain performance through succesive Procgen levels. In contrast, Pace exhibits robustness against this loss of plasticity, quickly achieving high performance with each new task.
We evaluate game environments including StarPilot, Dodgeball, Fruitbot, and Chaser. In all of these environments, we observe in Figure 4 that both Adam PPO and CReLU encounter a continually degrading loss of plasticity as these distribution shifts are introduced. In contrast, Pace PPO avoids this loss of plasticity, which contributes to its rapid reward increase when adapting to new levels. In the cumulative reward across all the Procgen levels, Pace PPO reveals normalized average improvements of 3,212.42% and 120.88% over Adam PPO and CReLU respectively (see Table 1). For later levels, in all games, Pace PPO’s reward does not decline as sharply as the baselines, potentially indicating forward transfer of skills from one level to the next.
Arcade Learning Environment (ALE) Atari
The Arcade Learning Environment (ALE) Atari 2600 benchmark is a collection of classic arcade games designed to assess reinforcement learning agents’ performance across a range of diverse gaming scenarios (Bellemare et al., 2012). We introduce distribution shifts by switching to a new Atari game every 4 million timesteps, where each game switch introduces a new task. This benchmark is more challenging compared to OpenAI Procgen: it requires the agent to handle distribution shifts in both the input (state) and the target (reward).
Figure 5: Reward in the lifelong Atari environments, across games with action spaces of 6 and 9. These plots demonstrate that Pace PPO rapidly adapts to new tasks, in contrast to the Adam PPO and CReLU which struggle to achieve high reward, indicating mild loss of plasticity.
In this experiment, we assessed two online settings distinguished by games with action spaces of 6 and 9. From Figure 5, both Adam PPO and CReLU sometimes failed to learn in certain games. In contrast, Pace PPO shows a substantial increase in reward over different games compared to the baselines. For example, during the first 12 million steps (3 games) in Atari 6, Pace PPO not only achieves a significantly higher mean reward but also demonstrates rapid reward increase. Over both experiment settings, Pace PPO shows an average normalized improvement of 329.73% over Adam PPO and 68.71% over CReLU (Table 1). In rare instances, such as the last 2 million steps of Atari 6, CReLU performs comparably to Pace PPO. This observation aligns with findings from (Abbas et al., 2023), which noted that CReLU tends to avoid plasticity loss in continual Atari setups.
Gym Control
We use the CartPole-v1 and Acrobot-v1 environments from the Gym Classic Control suite, along with LunarLander-v2 from Box2d Control. To introduce distribution shifts, a method outlined by Mendez et al. (2020) involves periodically altering the environment dynamics. Although such distribution shifts pose only mild challenges for robust methods like PPO with Adam (Appendix B). We instead implement a more challenging form of distribution shift. Every 200 steps we perturb each observation dimension with random noise within a range of ±$\pm 2$, treating each perturbation phase as a distinct task.
Figure 6: Reward performance across CartPole, Acrobot, and LunarLander Gym Control tasks. In these environments, both Adam PPO and CReLU experience extreme plasticity loss, failing to recover after the initial distribution shift. Conversely, Pace PPO successfully avoids such plasticity loss, rapidly adapting to each extreme distribution shift.
Here (Figure 6), we notice a peculiar behavior after introducing the first distribution shift in both Adam PPO and CReLU: policy collapse. We describe this as an extreme form of loss of plasticity. Surprisingly, Pace PPO remains resistant to these extreme distribution shifts. As we see in the Acrobot experiment, Pace PPO shows minimal to no policy damage after the first few distribution shifts, whereas Adam PPO and CReLU are unable to recover a policy at all. Across the three Gym Control environments, Pace PPO shows an average normalized improvement of 204.18% over Adam PPO and 1044.24% over CReLU (Table 1).
5 Discussion
Related work Combating loss of plasticity has been studied extensively in lifelong RL. A typical challenge for existing solutions is the tuning of their hyperparameters, which requires prior knowledge on the nature of the distribution shift, e.g., (Asadi et al., 2023; Ben-Iwhiwhu et al., 2023; Nikishin et al., 2023; Sokar et al., 2023; Mesbahi et al., 2024). An architectural modification called CReLU is studied in (Abbas et al., 2023), but our experiments suggest that its benefit might be specific to the Atari setup. Besides, Abel et al. (2018a, b) presented a theoretical analysis of skill transfer in lifelong RL, based on value iteration. Moreover, related contributions in nonstationary RL, where reward and state transition functions also change unpredictably, are limited to theoretical sequential decision-making settings with a focus on establishing complexity bounds (Roy et al., 2019; Cheung et al., 2020; Wei & Luo, 2021; Mao et al., 2022).
Our algorithm Pace builds on a long line of works on parameter-free OCO (see Section 3). To our knowledge, the only existing work applying parameter-free OCO to RL is (Jacobsen & Chan, 2021), which focuses on estimating of the value function (i.e. policy evaluation). Our scope is different, focusing on empirical RL in lifelong problems by exploring the key connection between parameter-free OCO and regularization.
Particularly, we are inspired by the Mechanic algorithm from (Cutkosky et al., 2023), which goes beyond the traditional convex setting of parameter-free OCO to handle stationary deep learning optimization tasks. Lifelong reinforcement learning, however, introduces a layer of complexity with its inherent nonstationarity. Furthermore, compared to Mechanic, Pace improves the scale tuner there (which is based on the coin-betting framework; Orabona & Pál, 2016) by the $\mathrm{erfi}$ algorithm that enjoys a better OCO performance guarantee. As an ablation study, we empirically compare Pace and Mechanic in the appendix (Figure 15). We find that Pace is slightly better, but both algorithms can mitigate the loss of plasticity, suggesting the effectiveness of the general “parameter-free” principle in lifelong RL.
On the choice of $\theta_{\mathrm{ref}}$
In general, the reference point $\theta_{\text{ref}}$ should be good or “safe” for Pace to perform effectively. One might presume that achieving this requires “warmstarting”, or pre-training using the underlying Base optimizer. While our experiments validate that such warmstarting is indeed beneficial (Appendix A), our main experiments show that even a random initialization of the policy’s weight serves as a good enough $\theta_{\text{ref}}$, even when tasks are similar (Figure 4).
This observation aligns with discussions by Lyle et al. (2023), Sokar et al. (2023), and Abbas et al. (2023), who suggested that persistent gradient steps away from a random initialization can deactivate ReLU activations, leading to activation collapse and loss of plasticity in neural networks. Our results also support Kumar et al. (2023)’s argument that maintaining some weights close to their initial values not only prevents dead ReLU units but also allows quick adaptation to new distribution shifts.
Tuning $L_{2}$ regularization
The success of Pace suggests that an adaptive form of regularization—anchoring to the reference point $\theta_{\mathrm{ref}}$—may suffice to counteract both mild and extreme forms of loss of plasticity. From this angle, we further elaborate the limitation of the $L_{2}$ regularization approach considered in (Kumar et al., 2023). It requires selecting a regularization strength parameter $\lambda$ through cross-validation, which is incompatible with the one-shot nature of lifelong learning settings. Furthermore, it is nontrivial to select the search grid: for example, we tried the $\lambda$-grid suggested by (Kumar et al., 2023), and there is no effective $\lambda$ value within the grid for the lifelong RL environments we consider. All the values are too small.
Continuing this reasoning, we conduct a hyperparameter search for $\lambda$, over various larger values $[0.2,0.8,1,5,10,15,20,25,30,35,40,45,50]$. Given the expense of such experiments, only the more sample-efficient control environments are considered. We discover that each environment and task responds uniquely to these regularization strengths (see bar plot of $\lambda$ values in Figure 7). This highlights the challenges of tuning $\lambda$ in a lifelong learning context, where adjusting for each environment, let alone each distribution shift, would require extensive pre-experimental analysis.
In contrast, Pace offers a parameter-free solution that adapts dynamically with the data in an online manner. The scaling output of Pace adjusts autonomously to the ongoing conditions, consistently competing with well-tuned $\lambda$ values in the various environments, as demonstrated in the reward plots for CartPole, Acrobot, and LunarLander (Figure 7).
Figure 7: For each Gym Control environment and the initial ten tasks, we identified the best $\lambda$, which is the regularization strength that maximizes reward for each task’s specific distribution shift. Additionally we determined the best overall (well-tuned) $\lambda$ for each environment, optimizing performance across all tasks (CartPole $\lambda=25$, Acrobot $\lambda=15$, LunarLander $\lambda=45$). The results demonstrate that each environment and each task’s distribution shift is sensitive to different $\lambda$ and that Pace PPO performs competitively with each environment’s well-tuned regularization strength.
Near convexity of lifelong RL
Our results demonstrate the rapid adaptation of Pace, in lifelong RL problems with complicated function approximation. From the perspective of optimization, the latter requires tackling both nonconvexity and nonstationarity, which is typically regarded intractable in theory. Perhaps surprisingly, when approaching this complex problem using the theoretical insights from OCO, we observe compelling results. This suggests a certain “hidden convexity” in this problem, which could be an exciting direction for both theoretical and empirical research (e.g., policy gradient methods provably converge to global optimizers in linear quadratic control (Hu et al., 2023)).
Limitations
While Pace offers robust adaptability in nonstationary environments, it can exhibit suboptimal performance at the outset due to the time it requires to “warm up”. In the early stages of deployment, Pace might underperform compared to the baseline optimizer. We address this by proposing a warmstarting solution detailed in Appendix A, which helps bridge the initial performance gap. Currently, our results demonstrate the effectiveness of Pace in conjunction with PPO. However, Pace is fundamentally an optimizer that could broadly apply to other RL algorithms. Future work should extend the application of Pace to additional algorithms such as Soft Actor-Critic (Haarnoja et al., 2018) and Deep Q-Networks (Mnih et al., 2013).
6 Conclusion
In this work, we introduced Pace, a parameter-free optimizer for lifelong RL that leverages the principles of OCO. Our approach dynamically refines regularization in a data-dependent manner, eliminating the need for hyperparameter tuning. Through extensive experimentation in Procgen, Atari, and Gym Control environments, we demonstrated that Pace effectively mitigates loss of plasticity and rapidly adapts to new distribution shifts, where baseline methods fail. Pace’s success leads to a compelling takeaway: empirical lifelong RL scenarios may exhibit more convex properties than previously appreciated, and might inherently benefit from parameter-free OCO approaches.
Therefore, in lifelong RL, to rapidly adapt to new changes without prior knowledge, we suggest you “Pick up (the) Pace”.
7 Acknowledgments
We thank Ashok Cutkosky for insightful discussions on online optimization in nonstationary settings. We are grateful to David Abel for his thoughtful insights on loss of plasticity in relation to lifelong reinforcement learning. We appreciate Kaiqing Zhang and Yang Hu for their comments on theoretical and nonstationary RL. This project is partially funded by Harvard University Dean’s Competitive Fund for Promising Scholarship.
Appendix
Appendix A Warmstarting
In our theoretical framework, we hypothesize that a robust parameter initialization, denoted as $\theta_{\mathrm{ref}}$, could enhance the performance of our models, suggesting that empirical implementations might benefit from initializing parameters using a base optimizer such as Adam prior to deploying Pace. Contrary to this assumption, our experimental results detailed in Section 4 reveal that warmstarting is not essential for Pace’s success. Below, we examine the performance of Adam PPO and Pace PPO when warmstarted.
Both Pace PPO and Adam PPO were warmstarted using Adam for the initial 150,000 steps in all games for the Atari and Procgen environments, and for the first 30 steps in the Gym Control experiments. As seen in Figure 8, in games like Starpilot, Fruitbot, and Dodgeball, Pace PPO surpasses Adam PPO in the first level/task of the online setup, with its performance closely matching that of Adam PPO in Chaser. Importantly, Pace PPO continues to circumvent the loss of plasticity encountered by Adam PPO, even when both are warmstarted. This makes sense since all of the distributions share some foundational game dynamics; the initial learning phases likely explore these dynamics, so leveraging a good parameter initialization to regularize in this early region can be beneficial for Pace—we observe that forward transfer occurs somewhat in later level distribution shifts as the reward does not drop back to zero where it initially started from.
Our findings indicate that warmstarting does not confer a significant advantage in the Atari games. This makes sense because a parameter initialization that is good in one game setting is likely a random parameterization for another setting, which is equivalent to the setup without warmstarting where Pace regularizes towards a random parameter initialization. In the Gym Control experiments although warmstarted Pace PPO manages to avoid the extreme plasticity loss and policy collapse seen in warmstarted Adam PPO, it does not perform as well as non-warmstarted Pace PPO. This variability underscores that the efficacy of warmstarting is environment-specific and highlights the challenge in predicting when Adam PPO may achieve a parameter initialization that is advantageous for Pace PPO to regularize towards.
From an overall perspective, warmstarting Pace PPO in every setting still shows substantial improvement over Adam PPO (Table 2)
Figure 8: Reward in the lifelong Procgen environments for StarPilot, Dodgeball, Fruitbot, and Chaser with warmstarted Pace PPO and warmstarted Adam PPO. Inital performance of Pace PPO is improved with warmstarting and continues to avoid loss of plasticity.
Figure 9: Reward in the lifelong Atari environments with warmstarted Pace PPO and warmstarted Adam PPO. No significant benefit is found by warmstarting Pace compared to not warmstarting it.
Figure 10: Reward in the lifelong Gym Control environments for CartPole-v1, Acrobot-v1, and LunarLander-v2 with warmstarted Pace PPO and warmstarted Adam PPO.
Table 2: Cumulative sum of mean episode reward over all distributions for Adam PPO warmstarted and Pace PPO warmstarted on Procgen, Atari, and Gym Control environments. Rewards are scaled by $10^{5}$; higher is better.
Environment Adam PPO Pace PPO (Ours)
Starpilot $3.0$ $\mathbf{10.2}$
Dodgeball $1.2$ $\mathbf{2.5}$
Chaser $1.3$ $\mathbf{1.6}$
Fruitbot $-0.4$ $\mathbf{0.6}$
CartPole $4.6$ $\mathbf{22.8}$
Acrobot $-142.9$ $\mathbf{-114.5}$
LunarLander $-190.7$ $\mathbf{-97.3}$
Atari6 $16.7$ $\mathbf{72.2}$
Atari9 $34.6$ $\mathbf{80.6}$
Appendix B Gravity Based Distribution Shifts
One method to introduce distribution changes in reinforcement learning environments is by altering the dynamics Mendez et al. (2020), such as adjusting the gravity in the CartPole environment. In this set of experiments, we manipulate the gravity by a magnitude of ten, randomly adding noise for one distribution shift, and then inversely, dividing by ten and adding random noise for the next shift. This process continues throughout the experiment.
Figure 11: Mean Episode Reward for Adam PPO on CartPole-v1 with varying gravity. Adam PPO demonstrates robust policy recovery across most gravity-based distribution shifts.
Our observations suggest that Adam PPO is robust to such dynamics-based distribution shifts, as shown in Figure 11. This indicates that while Adam PPO implicitly models the dynamics of the environment well—where changes in dynamics minimally impact performance—it struggles more with adapting to out-of-distribution observations such as seen in the main experiments (Figure 6) and in the warmstarting experiments (Figure 10).
Appendix C Scaling-Value Convergence
As discussed in the algorithm section (see Section 3), Pace operates as a meta-algorithm on top of a standard optimizer, denoted as Base. The crucial component of Pace involves the dynamic adjustment of the scaling parameter $S_{t+1}$, managed by the tuner algorithm (Algorithm 2). This parameter is data-dependent and typically ranges between $[0,1]$. The weight update $\theta_{t+1}$ is consequently defined as a convex combination of the current optimizer’s weight $\theta_{t}^{\textsc{Base}}$ and a predetermined reference point $\theta_{\text{ref}}$.
This section presents the convergence behavior of the scaling parameter $S_{t+1}$ across different environments, analyzed through the mean values over multiple seeds.
Figure 12: Convergence of the scaling parameter $S_{t+1}$ in the Procgen environments.
Figure 13: Evolution of the scaling parameter $S_{t+1}$ in the Atari environments. Here we don’t see a meaningful convergence of $S_{t+1}$.
Figure 14: Convergence of the scaling parameter $S_{t+1}$ in the Gym Control environments.
The convergence of the scaling parameter $S_{t+1}$ observed across the Procgen and Gym Control environments, as depicted in Figures 12 and 14, reflects a good scaling value that effectively determines the strength of regularization towards the initialization points, yielding robust empirical outcomes in lifelong RL settings. Interestingly, in Procgen environments, this converged scaling value exhibits consistency across various games, typically hovering between 0.02 and 0.03, as shown in Figure 12. In contrast, in the Gym Control environments, the scaling values are lower, ranging between 0.005 and 0.01, as illustrated in Figure 14.
Appendix D Comparison to Mechanic
In our analysis, we extend the examination to other OCO-based optimizers within the lifelong RL framework. Figure 15 presents a comparative assessment of Pace PPO and Mechanic PPO (Cutkosky et al., 2023) in the Starpilot game from the Procgen suite. The results indicate that while Mechanic effectively mitigates plasticity loss and adapts quickly to new distribution shifts, it slightly underperforms in comparison to Pace. This suggests potential for further explorations and a broader evaluation of OCO-based optimizers in lifelong learning environments.
Figure 15: Reward in the lifelong StarPilot environment with both Pace PPO and Mechanic PPO. Mechanic PPO performs similarly to Pace PPO, although slightly underperforming Pace PPO.
Appendix E Experimental Setup
Procgen and Atari Vision backbone
For both the Atari and Procgen experiments, the Impala architecture was used as the vision backbone. The Impala model had 3 Impala blocks, each containing a convolutional layer followed by 2 residual blocks. The output of this is flattened and connected to a fully connected layer. The impala model parameters are initialized using Xavier uniform initialization.
Policy and Value Networks
Across all experiments—including Gym Control, Atari, and Procgen—the policy and value functions are implemented using a multi-layer perceptron (MLP) architecture. This architecture processes the input features into action probabilities and state value estimates. The MLP comprises several fully connected layers activated by ReLU. The output from the final layer uses a softmax activation.
Table 3: PPO Parameters for Atari, Procgen, and Gym Control Experiments
Parameter Atari Procgen Control
Steps per update 2,000 1,000 800 (2 episodes with 400 steps)
Batch size 250 125 32
Epochs per update 3 3 5
Epsilon clip for PPO 0.2 0.2 0.2
Value coefficient 0.5 0.5 0.5
Entropy coefficient 0.01 0.01 0.01
Base Optimizer Adam (LR: 0.001) Adam (LR: 0.001) Adam (LR: 0.01)
Architecture Impala + MLP Impala + MLP MLP
Pace
Pace, for all experiments, was implemented using the same experiment-specific baseline architectures and baseline optimizer. For the Procgen and Atari experiments, the base Adam optimizer was configured as the same as baseline, with a learning rate of 0.001, and for the Gym Control experiments, a learning rate of 0.01 was used. Both learning rates were tested for all experiments and found to have negligible differences in performance outcomes. Other than the learning rate, we use the default Adam parameters, including weight decay and betas, followed by the specifications outlined in the PyTorch Documentation. [^4] https://pytorch.org/docs/stable/generated/torch.optim.Adam.html
The setup for Pace included $\beta$ values for adaptive gradient adjustments: 0.9, 0.99, 0.999, 0.9999, 0.99999, and 0.999999. Both $S_{t}$ and $\varepsilon$ were initially set to ($1\times 10^{-8}$). Modifications were made to a PyTorch error function library, which accepts complex inputs to accommodate the necessary computations for the imaginary error function. This library can be found at Torch Erf GitHub. [^5] https://github.com/redsnic/torch_erf
Distribution Shifts
In the Atari experiments, game environments were switched every 4 million steps. The sequence for games with an action space of 6 included “BasicMath”, “Qbert”, “SpaceInvaders”, “UpNDown”, “Galaxian”, “Bowling”, “Demonattack”, “NameThisGame”, while games with an action space of 9 included “LostLuggage”, “VideoPinball”, “BeamRider”, “Asterix”, “Enduro”, “CrazyClimber”, “MsPacman”, “Koolaid”.
For Procgen experiments, individual game levels were sampled using a seed value as the start_level parameter, which was incremented sequentially to generate new levels. Each new environment was introduced every 2 million steps.
In the Gym Control experiments, each observation dimension was randomly perturbed by a value ranging from 0 to 2. This perturbation was constant for 200 timesteps, after which a new perturbation was applied, effectively switching the environmental conditions every 200 steps.
Statistical Significance
The Procgen and Atari experiments were conducted with 8 seeds/runs, while the Gym Control experiments utilized 25 seeds/runs. The exception was in the $L_{2}$ initialization experiments, which used 15 seeds/runs per regularization strength. In Figures 4, 5, 6, 7, 8, 9, 10, 11, the plotted lines represent the mean of all of the mean episode rewards from the different seeds/runs, and the shaded error bands indicate the standard deviation of all of the mean episode rewards from the different seeds/runs.
Compute Resources
For the Procgen and Atari experiments, each was allocated a single A100 GPU, typically running for 3-4 days to complete. The Gym Control experiments were conducted using dual-core CPUs, generally concluding within a few hours. In both scenarios, an allocation of 8GB of RAM was sufficient to meet the computational demands.